4. A pulley 600 mm. in diameter transmits 40 kW at 500 rpm. The arc of contact between the belt and pulley is 144° having a center distance of 1100 mm. Determine: (a) diameter of the driven pulley for open belt drive, (b) speed of the driven pulley, and (c) belt length for open belt connection.


1
Expert's answer
2022-03-19T02:46:19-0400

"P =\frac{2\pi NT}{60}"

"T =\frac{40000\times 60}{2\pi\times500 } = 763.94" Nm

(a) Diameter of the driven pulley for open belt drive

"sin \beta = \frac{D-d}{2C}"

"D = Sin \beta \times 2C +d =sin 144 \times 2200 + 600 = 1893.13" mm

(b) Speed of the driven pulley

"\frac{n}{N}=\frac{D}{d}"


"n = N\times \frac{D}{d}= 500\times \frac{1893.13}{600}= 1577.61 \approx 1578" rpm


(c) belt length for open belt connection


"L = 2C+\frac{\pi(D+d)}{2}+\frac{(D-d)^2}{4C} =2\times 1100+\frac{\pi (1893.13+600)}{2}+\frac{(1893.13-600)^2}{4\times1100}=6496.24" mm